The no. of degrees in one acute angle of a right-angled triangle is equal to the no. of grades in the other acute angle of the same triangle. Express both the angles in degrees.
Please give the solution.

Grade:11

1 Answers

sunil raikwar
10 years ago
we Know that, 180 degree = 200 grade therefore 1 grade = 9/10 degree let x degree + x grade =90 degree x degree + 9x/10 degree = 90 degree 19x/10 = 90 degree x = 900/19 so angles are 900/19 degree & 810/19 degree